Equivalent weight of H3PO2(molecular weight =m ) when it disproportionates into PH3 and H3PO3
M
M/2
M/4
3M/4

Grade:10

1 Answers

Rituraj Tiwari
3 years ago
For disproportionation reactions, use a trick to find equivalent mass; E = E1 + E2; where E1 and E2 are equivalent masses of oxidation and reduction half reactions of same element.

Now, find the n-factor(total change in oxidation number per molecule) and then equivalent mass = molar mass/n-factor.

So, when H3PO2 changes into PH3, oxidation state of phosphorous changes from +1 to -3 so as n-factor is 4. Also when H3PO2 changes into H3PO4, oxidation state of phosphorous changes from +1 to +5 so as n-factor is again 4.

Now equivalent mass of H3PO2, E =(M/4) + (M/4) = M/2